OpenStudy (anonymous):

x=2 2x-y=4 3x+2y-2z=4 solve by substitution

OpenStudy (anonymous):

If x=2, put x into the second equation. 2(2)-y=4 4-y=4 -y=0 y=0. Now you know y. Plug x and y into the last equation. 3(2)+2(0)-2z=4 6+0-2z=4 -2z=-2 z=1. x=2, y=0, z=1.

OpenStudy (anonymous):

oh ok thanks...

OpenStudy (anonymous):

Yep. :)